OpenStudy (anonymous):

HELP PLS!!! 2C4H10(g)+13O2(g)→10H2O(g)+8CO2(g) Calculate the mass of water produced when 7.36g of butane reacts with excess oxygen. Express your answer to three significant figures and include the appropriate units. Calculate the mass of butane needed to produce 72.7g of carbon dioxide. Express your answer to three significant figures and include the appropriate units. how do you do it?

OpenStudy (aaronq):

First convert the mass values to moles (n), moles (n)=mass (m)/Molar mass(M) equation (1) then use the stoichiometric coefficients to build a ratio: ------------------------------------------------------------------------- For a general equation: aA +bB -> cC where a,b,c (lowercase) are the coefficients and A,B,C are the species \[\frac{ n _{A} }{ a }=\frac{ n _{B} }{ b }=\frac{ n _{C}}{ c }\] this way you can "normalize" the amount of moles for the reaction, (you can find the amount of moles so they're reacting in a 1 :1 ratio). -------------------------------------------------------------------------- so for example, 2C4H10(g)+13O2(g)→10H2O(g)+8CO2(g) Calculate the mass of water produced when 7.36g of butane reacts with excess oxygen. once you convert the mass of butane to moles, write their relationship like this (notice each one is divided by it's own coefficient) \[\frac{ n _{butane} }{ 2 }=\frac{ n _{water} }{ 10}\] then plug in moles of butane, and solve for moles of water for example if moles of butane =5 then 5/2=n/10 , n = 25 moles of water (this is NOT the answer but an example) they're asking for the mass, so convert to back to mass using equation (1).